OpenStudy (anonymous):

What is the vertex of the graph of y = −1(x − 2) ^2 + 3? Someone help please explain it to me.

OpenStudy (lochana):

y = -(x-2)^2 +3 = (-x^2 +4x -4) + 3 therefor y = - ( x^2 -4x+ 1) then y = -{ (x-2)^2 -4 +1} = -{(x-2)^2 - 3} y= 3 - (x-2)^2 now we can say if (x-2)^2=0 and then y become its highest value.which is 3, (i'm not quite sure)

OpenStudy (anonymous):

it is supposed to be a pair (3, −2) (−2, 3) (3, 2) (2, 3)

OpenStudy (lochana):

i think it's (2,3)

OpenStudy (anonymous):

okay thank you